### v1

@@ E.2 Emotions: Regression Tables
Table 9: Emotions: Reduced Exposure Experiment, In-Feed. Regression tables of the analysis of the emotion outcomes for the in-feed survey responses in the reduced exposure experiment.
@@
Condition: Treatment $5.13^{*}$ $4.38^{*}$ $-0.09$ $-2.69$
$[2.05;8.22]$ $[1.00;7.76]$ $[-3.75;3.57]$ $[-6.09;0.70]$
Baseline: Mean Answer $0.56^{*}$ $0.55^{*}$ $0.68^{*}$ $0.66^{*}$
@@
Recruitment Platform: CloudResearch $1.46$ $5.10^{*}$ $-6.31^{*}$ $-3.93^{*}$
$[-1.65;4.57]$ $[1.70;8.50]$ $[-9.99;-2.62]$ $[-7.35;-0.51]$
@@
Table 10: Emotions: Increased Exposure Experiment, In-Feed. Regression tables of the analysis of the emotion outcomes for the in-feed survey responses in the increased exposure experiment.
@@
Recruitment Platform: CloudResearch $0.06$ $0.05$ $-0.07$ $-0.04$
$[-0.08;0.21]$ $[-0.08;0.18]$ $[-0.22;0.08]$ $[-0.20;0.11]$
@@
Table 11: Emotions: Reduced Exposure Experiment, Post-Experiment. Regression tables of the analysis of the emotion outcomes for the post-experiment survey responses in the reduced exposure experiment.
@@
Recruitment Platform: CloudResearch $0.08$ $0.17^{*}$ $-0.25^{*}$ $-0.12$
$[-0.08;0.24]$ $[0.01;0.34]$ $[-0.42;-0.08]$ $[-0.28;0.05]$
@@
Table 12: Emotions: Increased Exposure Experiment, Post-Experiment. Regression tables of the analysis of the emotion outcomes for the post-experiment survey responses in the increased exposure experiment.
@@
Recruitment Platform: CloudResearch $0.62$ $2.52$ $-4.23^{*}$ $-3.10$
$[-2.07;3.32]$ $[-0.26;5.30]$ $[-7.71;-0.76]$ $[-6.24;0.04]$
@@
Table 13: Full population – Emotions: Reduced Exposure Experiment, In-Feed. Regression tables of the analysis of the emotion outcomes for the in-feed survey responses in the reduced exposure experiment.
@@
Recruitment Platform: CloudResearch $1.95$ $5.04^{*}$ $-6.02^{*}$ $-3.70^{*}$
$[-1.10;5.00]$ $[1.70;8.37]$ $[-9.65;-2.40]$ $[-7.05;-0.35]$
@@
Table 14: Full population – Emotions: Increased Exposure Experiment, In-Feed. Regression tables of the analysis of the emotion outcomes for the in-feed survey responses in the increased exposure experiment.
@@
Condition: Treatment $-0.08$ $-0.00$ $0.04$ $0.04$
$[-0.21;0.04]$ $[-0.12;0.12]$ $[-0.09;0.18]$ $[-0.10;0.17]$
Pre-survey Answer $0.45^{*}$ $0.44^{*}$ $0.54^{*}$ $0.41^{*}$
@@
Recruitment Platform: CloudResearch $0.11$ $0.09$ $-0.08$ $-0.04$
$[-0.03;0.24]$ $[-0.04;0.21]$ $[-0.22;0.06]$ $[-0.18;0.11]$
@@
Table 15: Full population – Emotions: Reduced Exposure Experiment, Post-Experiment. Regression tables of the analysis of the emotion outcomes for the post-experiment survey responses in the reduced exposure experiment.
@@
Condition: Treatment $-0.02$ $0.08$ $0.09$ $-0.07$
$[-0.17;0.13]$ $[-0.07;0.23]$ $[-0.06;0.25]$ $[-0.23;0.09]$
Pre-survey Answer $0.58^{*}$ $0.55^{*}$ $0.56^{*}$ $0.43^{*}$
@@
Recruitment Platform: CloudResearch $0.13$ $0.22^{*}$ $-0.27^{*}$ $-0.12$
$[-0.03;0.28]$ $[0.06;0.37]$ $[-0.43;-0.11]$ $[-0.28;0.05]$
@@
Table 16: Full population – Emotions: Increased Exposure Experiment, Post-Experiment. Regression tables of the analysis of the emotion outcomes for the post-experiment survey responses in the increased exposure experiment.
E.3 Political attitude shifts
As preregistered, we report the effect of the intervention on the seven remaining attitudes we used to identify the AAPA posts. We find that Reduced Exposure to AAPA content reduces the Biased Evaluation of Politicized Facts while increasing the exposure increases the Opposition to Bipartisan Cooperation (Fig. 16). However, none of these effects are significant after adjusting for multiple hypotheses testing as preregistered (Sec. D).
Figure 16: Treatment effects on anti-democratic attitudes estimated by regressing the participants’ post-survey responses on the treatment indicator, their pre-survey response to the same question, and an indicator of the platform they were recruited on. The p values are adjusted for multiple hypotheses testing as preregistered.
@@ E.3 Political attitude shifts
Figure 17: Each value represents the estimated effect from a separate regression model. The dependent variable is the political feeling response in the post-experiment survey, controlling for both the pre-experiment survey response and the fraction of political posts viewed over a ten-day period for at least one second, which expressed support for a particular factor (i.e., $post\_experiment\_rating\sim pre\_experiment\_rating+fraction\_v_{n}$, where $v_{n}$ is a specific factor such as partisan animosity or social distrust). The coefficient represents the estimated change (on a 0-100 scale) in feelings toward people from the opposite party when 100% of the viewed political posts expressed support for the corresponding factor ($v_{n}$). All participants are included in the analysis (N=1,256).
E.4 Engagement
During the baseline period, we observed that political posts enjoyed a higher reshare rate: 0.677%, compared to 0.484% for non-political posts (a relative increase of 39.7%; U=714,022; P$=$0.046). Similarly, political content had a higher favorite rate, with an average engagement rate for posts viewed for at least one second of 5.07% vs. 4.29% for non-political posts (U=645,266; P$=$0.001). We observed that AAPA posts, while shared more frequently than non-AAPA content, did not exceed political non-AAPA posts in share rate (0.662% vs. 0.674%; U=668,945, P=0.003). Similarly, although AAPA content had a higher favorite rate than non-AAPA posts, it was favorited less than non-AAPA political posts (4.6% vs. 5.23%; U=711,396, P$<$0.001).
When examining the effects of our interventions, none of the differences reached statistical significance. In the Increased Exposure experiment, the average number of daily sessions per participant–defined as scrolling activities separated by at least one hour–was 1.65 in the control group and 1.53 in the treatment group (U=32,220; P=0.18). In the Reduced Exposure experiment, the control and treatment groups averaged 1.53 and 1.49 sessions per day, respectively (U=66,940; P=0.71). After controlling for the average daily sessions during the baseline period and the recruitment platform, the change in session count was not statistically significant in either experiment (Increased Exposure: -0.07 sessions, CI [-0.19, 0.04]; P=0.197; 95% Reduced Exposure: -0.04 sessions, CI [-0.13, 0.05] P=0.381; 95%). The differences in the time spent on the platform across conditions also do not reach statistical significance. In the Increased Exposure experiment, participants in the control condition spent on average of 44.4 minutes vs. 41.3 minutes in the treatment condition (U=37,545; P=0.12), while in the Reduced Exposure, they spent 46.8 minutes in the control and 42.2 minutes in the treatment condition (U=70,678; P=0.10).
Similarly, repost rates during the experimental period did not differ significantly between treatment and control groups. In the Increased Exposure experiment, the repost rate was 0.52% in the control group and 1.07% in the treatment group (U=30,859; P=0.83). In the Reduced Exposure experiment, the repost rates were 0.74% for the control group and 0.46% for the treatment group. When controlling for user baseline averages and the platform using regression analysis (Fig. 18), the changes in repost rates remained non-significant (Increased Exposure: increase of 0.27%, 95% CI [-0.002, 0.008]; P=0.29; Reduced Exposure: reduction of -0.06%, 95% CI [-0.002, 0.001] P=0.503;).
The favorite rates during the experimental period also showed no significant differences. The overall favorite rate during the baseline was 4.73%. Controlling for user baseline and platform (Fig. 19), the favorite rate changed by 0.14% in the Increased Exposure group (95% CI [-0.009, 0.012]; P=0.796) and decreased by 0.29% in the Reduced Exposure group (95% CI [-0.010, 0.004]; P=0.43).
Finally, we observed no significant differences between treatment and control groups in terms of the number of new posts or replies during the experimental period.
Figure 18: Treatment effect on repost rate. Coefficient of a linear regression predicting participants’ repost rate during the experiment based on their baseline repost rate and experimental condition (treatment or control). Error bars represent 95% confidence intervals.
Figure 19: Treatment effect on favorite rate. Coefficient of a linear regression predicting participants’ favorite rate during the experiment based on their baseline favorite rate and experimental condition (treatment or control). Error bars represent 95% confidence intervals.
E.5 Heterogeneous Treatment Effects
Fig. 20 and Fig. 21 show the heterogeneous treatment effects on affective polarization in the two experiments. We use the same regression specifications as in the main analyses but interact the treatment with the moderator. As in the covariate balance and attrition tests, we coarsened the education, income, and SES ladder covariates to avoid small categories. We consider only a subset of covariates and split them into primary and secondary moderators as preregistered. After adjusting the p-values for multiple hypotheses testing, none of the covariates reach statistical significance.
Figure 20: Heterogeneous treatment effect estimates of reducing and increasing exposure to AAPA content on affective polarization based on the in-feed measurements. The p values are adjusted for multiple hypotheses testing as preregistered.
Figure 21: Heterogeneous treatment effect estimates of reducing and increasing exposure to AAPA content on affective polarization based on the post-experiment measurements. The p values are adjusted for multiple hypotheses testing as preregistered.
E.6 Qualitative feedback on the impact of the experiment
@@
At the end of the study, in the post-survey, we asked participants: “Did the Chrome extension affect your experience on Twitter/X in any way?” This question was open-ended, and 1,218 participants (96.9%) provided a non-empty response. Most answers were short and indicated no impact, such as “No” or “Not that I could tell.” We individually coded each response with the support of AI (GPT-4), providing the original survey question for context and asking whether the participant reported any impact on their experience on Twitter/X.
@@
To investigate this further, we adopted a quantitative approach by embedding participant responses using a pre-trained BERT-based model (all-MiniLM-L6-v2). We then predicted the assigned experiment using logistic regression with L2 regularization, employing 5-fold cross-validation. The classification achieved an F1 score of 53.9% ($\pm$ 4.7%) on a balanced dataset. Similarly, we predicted the experimental condition (treatment or control) for each experiment. For the reduced experiment, the F1 score was 47.4% ($\pm$ 6.8%), and for the increased experiment, it was 50.5% ($\pm$ 4.7%).
@@
These results suggest that the experimental condition did not significantly impact participants’ experience, as they did not perceive or report notable differences.

### v2

@@ Acknowledgments
Table S23: Emotions: Reduced Exposure Experiment, In-Feed. Regression tables of the analysis of the emotion outcomes for the in-feed survey responses in the reduced exposure experiment.
@@
Condition: Treatment $5.13^{*}$ $4.38^{*}$ 0.09-0.09 2.69-2.69
$[2.05;8.22]$ $[1.00;7.76]$ [3.75;3.57][-3.75;3.57] [6.09;0.70][-6.09;0.70]
Baseline: Mean Answer $6^{*} 0.550$$5^{*}$ $8^{*} 0.660$$6^{*}$
@@
Recruitment Platform: CloudResearch $1.46$ $5.10^{*}$ $6.31^{*}$ $3.93^{*}$
[1.65;4.57][-1.65;4.57] $[1.70;8.50]$ [9.99;2.62][-9.99;-2.62] [7.35;0.51][-7.35;-0.51]
@@
Table S24: Emotions: Increased Exposure Experiment, In-Feed. Regression tables of the analysis of the emotion outcomes for the in-feed survey responses in the increased exposure experiment.
@@
Recruitment Platform: CloudResearch $0.06$ $0.05$ 0.07-0.07 0.04-0.04
[0.08;0.21][-0.08;0.21] [0.08;0.18][-0.08;0.18] [0.22;0.08][-0.22;0.08] [0.20;0.11][-0.20;0.11]
@@
Table S25: Emotions: Reduced Exposure Experiment, Post-Experiment. Regression tables of the analysis of the emotion outcomes for the post-experiment survey responses.
@@
Recruitment Platform: CloudResearch $0.08$ $0.17^{*}$ $0.25^{*}$ 0.12-0.12
[0.08;0.24][-0.08;0.24] $[0.01;0.34]$ [0.42;0.08][-0.42;-0.08] [0.28;0.05][-0.28;0.05]
@@
Table S26: Emotions: Increased Exposure Experiment, Post-Experiment. Regression tables of the analysis of the emotion outcomes for the post-experiment survey responses.
@@
Recruitment Platform: CloudResearch $0.62$ $2.52$ $4.23^{*}$ 3.10-3.10
[2.07;3.32][-2.07;3.32] [0.26;5.30][-0.26;5.30] [7.71;0.76][-7.71;-0.76] [6.24;0.04][-6.24;0.04]
@@
Table S27: Full population – Emotions: Reduced Exposure Experiment, In-Feed. Regression tables of the analysis of the emotion outcomes for the in-feed survey responses.
@@
Recruitment Platform: CloudResearch $1.95$ $5.04^{*}$ $6.02^{*}$ $3.70^{*}$
[1.10;5.00][-1.10;5.00] $[1.70;8.37]$ [9.65;2.40][-9.65;-2.40] [7.05;0.35][-7.05;-0.35]
@@
Table S28: Full population – Emotions: Increased Exposure Experiment, In-Feed. Regression tables of the analysis of the emotion outcomes for the in-feed survey responses.
@@
Condition: Treatment 0.08-0.08 0.00-0.00 $0.04$ $0.04$
[0.21;0.04][-0.21;0.04] [0.12;0.12][-0.12;0.12] [0.09;0.18][-0.09;0.18] [0.10;0.17][-0.10;0.17]
Pre-survey Answer $5^{*} 0.440$$4^{*}$ $0.54^{*}$ $0.41^{*}$
@@
Recruitment Platform: CloudResearch $0.11$ $0.09$ 0.08-0.08 0.04-0.04
[0.03;0.24][-0.03;0.24] [0.04;0.21][-0.04;0.21] [0.22;0.06][-0.22;0.06] [0.18;0.11][-0.18;0.11]
@@
Table S29: Full population – Emotions: Reduced Exposure Experiment, Post-Experiment. Regression tables of the analysis of the emotion outcomes for the post-experiment survey responses.
@@
Condition: Treatment 0.02-0.02 $0.08$ $0.09$ 0.07-0.07
[0.17;0.13][-0.17;0.13] [0.07;0.23][-0.07;0.23] [0.06;0.25][-0.06;0.25] [0.23;0.09][-0.23;0.09]
Pre-survey Answer $8^{*} 0.550$$5^{*}$ $0.56^{*}$ $0.43^{*}$
@@
Recruitment Platform: CloudResearch $0.13$ $0.22^{*}$ $0.27^{*}$ 0.12-0.12
[0.03;0.28][-0.03;0.28] $[0.06;0.37]$ [0.43;0.11][-0.43;-0.11] [0.28;0.05][-0.28;0.05]
@@
Table S30: Full population – Emotions: Increased Exposure Experiment, Post-Experiment. Regression tables of the analysis of the emotion outcomes for the post-experiment survey responses.
@@
Figure S22: Relationship between the effect on affective polarization (post-experiment) and exposure to posts with each of the eight AAPA factors. Each value represents the estimated effect from a separate regression model. The dependent variable is the political feeling response in the post-experiment survey, controlling for both the pre-experiment survey response and the fraction of political posts viewed over a ten-day period for at least one second, which expressed support for a particular factor (i.e., $\text{{affective\_polarization}}_{\text{{post}}}\sim\text{{affective\_polarization}}_{\text{{pre}}}+\text{{fraction\_}}v_{n}$, where $v_{n}$ is a specific factor such as partisan animosity or social distrust). The coefficient represents the estimated change (on a 0-100 scale) in feelings toward people from the opposite party when 100% of the viewed political posts expressed support for the corresponding factor ($v_{n}$). All participants who completed the post-experiment survey are included in the analysis (N = 1,256).
S11.3 AAPA Factors Analysis
To investigate how individual AAPA factors influence in-feed survey responses, we first analyzed the relationship between participants’ reported out-party feelings and the aggregate AAPA score of the post immediately preceding each in-feed survey. Using a mixed-effects regression model with participants as a random effect, we found a significant negative association (coefficient = –0.124; 95% CI: [–0.236, –0.012]; P = 0.029). This suggests that posts containing more AAPA factors were associated with colder feelings toward the opposing party.
We then extended this analysis to test the effect of individual AAPA factors, included as binary indicators in the same mixed-effects framework. However, none of the individual factors reached statistical significance, likely due to limited statistical power and collinearity among the predictors.
To further examine the impact of these factors, we investigated the correlation between the participants’ response to the post-experiment out-party feeling thermometer and the overall exposure to posts with a certain factor. As shown in Fig. S22, certain types of AAPA factors, such as expressions of partisan animosity, opposition to bipartisan collaboration, and biased evaluations of political facts, do predict changes in participants’ post-experiment attitudes. This figure reports the variation in post-experiment scores as a function of the fraction of content consumed that reflects each specific factor. For example, the regression coefficients suggest that if a participant’s feed contained exclusively (100%) posts expressing opposition to bipartisan collaboration, their attitude would shift by approximately 15 degrees. It is essential to note that other types of content, such as support for political violence, contribute to the final score and may influence changes in attitude. However, their effects may not reach statistical significance, possibly because they are rare (appearing in only 5.3% of political posts and 1.62% of the posts seen by participants in their feeds), and we are not powered to detect effects for these rarer factors when analyzed individually.
S11.4 Qualitative Feedback on the Impact of the Experiment
@@
At the end of the study, in the post-experiment survey, we asked participants: “Did the Chrome extension affect your experience on Twitter/X in any way?” This question was open-ended, and 1,218 participants (96.9%) provided a non-empty response. Most answers were short and indicated no impact, such as “No” or “Not that I could tell.” We individually coded each response with the support of AI (GPT-4), providing the original survey question for context and asking whether the participant reported any impact on their experience on Twitter/X.
@@
To investigate this further, we adopted a quantitative approach by embedding participant responses using a pre-trained BERT-based model (all-MiniLM-L6-v2). We then predicted the assigned experiment using logistic regression with L2 regularization, employing 5-fold cross-validation. The classification achieved an F1 score of 53.9% ($\pm$ 4.7%) on a balanced dataset, performance statistically indistinguishable from random guessing. Similarly, we predicted the experimental condition (treatment or control) for each experiment. For the reduced experiment, the F1 score was 47.4% ($\pm$ 6.8%), and for the increased experiment, it was 50.5% ($\pm$ 4.7%); performance no better than random guessing.
@@
Together, these results suggest that the experimental condition did not significantly impact participants’ experience, as they did not perceive or report notable differences.
S11.5 Decay of Treatment Effects
To investigate whether there is any short-term decay in the treatment effects of the interventions reducing and increasing exposure to AAPA content, we examined (1) how the treatment effects varied over the course of the experiment and (2) whether the time gap between the end of the intervention and participants’ completion of the post-experiment survey was correlated with the size of the treatment effects.
To examine how the treatment effects varied as the experiment progressed, we reanalyzed the in-feed responses, extending our model specification to include an indicator variable for the day of the experiment (relative to the start of the intervention for each participant) and its interaction with the experimental condition indicator. Figure S23 shows the point estimates and 95% confidence intervals of the interaction coefficients. We find that none of the interactions between the experimental condition and day of the experiment are statistically significant in either the Reduced or Increased Exposure experiments. These results suggest that the treatment effects do not significantly vary over time and that there is no evidence of decay effects within the week of the intervention.
Figure S23: Variation in treatment effects on affective polarization for each day of the intervention, relative to the first day, as measured by the in-feed surveys. Points show the coefficients of the interactions between treatment assignment and experiment day, and error bars represent the 95% CIs.
We also tested for potential decay effects based on the timing of participants’ last exposure to treatment. We can perform this analysis because of natural variation in the time delta between (a) when our browser extension automatically turned off the reranking at the end of the final day of the experiment and (b) when participants filled out our post-survey. Participants could take the survey immediately after the reranking turned off (after midnight on the last day of the experiment), but most of them completed the survey in the following days. In the interim the participants could still browse their (unmodified) X feeds, but were shown a large banner with a link to the post-experiment survey. We extended our model specification for analyzing the post-experiment effects to include the total hours between when a participant last viewed a post (for at least one second) during the experiment and when they began the post-survey. The median delay gap was 25.5 hours (mean = 42.3). Table S31 shows the results of the regression analysis. We find that the time difference between the last exposure and the start of the post-survey is not a significant predictor of changes in affective polarization and does not alter the significance of the treatment effect.
Together, these results suggest that the treatment effects do not decay in the short term.
Post-Experiment
Reduced Exposure Increased Exposure
(Intercept) $0.41$ $4.59^{*}$
[1.88;2.70][-1.88;2.70] $[1.70;7.49]$
Condition: Treatment $2.10^{*}$ $2.44^{*}$
$[0.14;4.05]$ [4.76;0.13][-4.76;-0.13]
Pre-survey Answer $0.81^{*}$ $0.83^{*}$
$[0.77;0.85]$ $[0.77;0.88]$
Recruitment Platform: CloudResearch 0.54-0.54 $3.04^{*}$
[2.60;1.52][-2.60;1.52] [5.50;0.57][-5.50;-0.57]
Hours Difference $0.02$ 0.02-0.02
[0.01;0.05][-0.01;0.05] [0.05;0.02][-0.05;0.02]
Num. obs. $622$ $468$
Null hypothesis value outside the confidence interval.
Table S31: Affective Polarization: Post-Experiment. Regression tables of the analysis of the affective polarization outcome for the post-experiment survey responses, including the number of hours from the last post view (at least 1 second) and the answer to the post-experiment survey.
Figure S24: Standard deviation of daily participant responses (difference from baseline) by day, experiment, and condition. Shaded areas represent 95% bootstrapped confidence intervals.
S11.6 Survey Fatigue and Response Lock-in
A potential risk in our experimental design is that participants repeatedly asked to report their feelings toward political outgroups may experience survey fatigue and settle on a fixed response over time, responding habitually rather than reflectively. This may cause responses to converge to a preferred value, regardless of the actual stimuli. To examine this possibility, we computed, for each day, the standard deviation of the difference between each participant’s in-feed response and their own baseline. This metric captures the dispersion of daily responses relative to each participant’s initial reporting level, aggregated across users. If participants were anchoring to a fixed response, we would expect this marginal standard deviation to decline over time. However, as shown in Figure S24, we observe no consistent downward trend in either experiment or condition. This suggests that participants continued to respond with a similar degree of variation, providing no evidence of convergence.
S11.7 Controlling for Major Political Events
The experiment was conducted during a period of heightened political activity preceding the 2024 US election. Several major events during this time may have influenced participants’ feelings toward the opposing party, potentially affecting levels of affective polarization. These events include:
  • The assassination attempt on Donald Trump (July 13),
  • The Republican National Convention (July 15–18),
  • Joe Biden dropping out of the presidential race (July 21),
  • The presumptive nomination of Kamala Harris (July 22),
  • Kamala Harris’s official nomination as the Democratic candidate (August 5).
To assess whether these events were driving the observed effects, we conducted an additional analysis using binary indicator variables. Each indicator was set to 1 if a survey response occurred on the day of a given event or the day immediately following it. The exception was the Republican National Convention, for which the indicator covered the entire three-day span. These event indicators were then interacted with the treatment variable in our regression models. The results, summarized in Table S32, show that our core findings remain robust: the treatment variable remains statistically significant even after accounting for these events.
In the Reduced Exposure experiment, none of the event indicators had a statistically significant interaction with the treatment, indicating that the observed effects are not attributable to these external events. In the Increased Exposure experiment, the main treatment effect remains statistically significant. While some event indicators—such as the assassination attempt on Trump and the presumptive nomination of Harris—reach the conventional 5% significance threshold, these effects do not survive Bonferroni correction for multiple comparisons (five hypotheses). Furthermore, their interaction terms with the treatment condition are not significant after correction.
Overall, these results suggest that the main effects of our intervention on affective polarization persist even when accounting for the influence of major political events during the study period.
In-Feed
Reduced Exposure Increased Exposure
(Intercept) 0.14-0.14 $3.12^{*}$
[3.01; 2.73][-3.01;\ 2.73] [1.11; 5.14][1.11;\ 5.14]
Recruitment Platform: CloudResearch 1.32-1.32 1.05-1.05
[4.15; 1.51][-4.15;\ 1.51] [3.09; 0.99][-3.09;\ 0.99]
Treatment $3.88^{*}$ $2.65^{*}$
[1.14; 6.61][1.14;\ 6.61] [4.68;0.62][-4.68;\ -0.62]
Baseline: Outparty Feeling $0.92^{*}$ $0.84^{*}$
[0.85; 1.00][0.85;\ 1.00] [0.80; 0.88][0.80;\ 0.88]
Assassination Attempt 1.72-1.72 $3.77^{*}$
[4.62; 1.17][-4.62;\ 1.17] [6.63;0.91][-6.63;\ -0.91]
Assassination Attempt $\times$ Treatment $1.64$ $4.50^{*}$
[2.67; 5.95][-2.67;\ 5.95] [0.41; 8.58][0.41;\ 8.58]
Republican Convention 0.37-0.37 0.45-0.45
[2.35; 1.61][-2.35;\ 1.61] [2.49; 1.59][-2.49;\ 1.59]
Republican Convention $\times$ Treatment 0.20-0.20 1.87-1.87
[3.08; 2.68][-3.08;\ 2.68] [4.93; 1.20][-4.93;\ 1.20]
Biden Dropout $0.01$ 0.48-0.48
[1.33; 1.35][-1.33;\ 1.35] [1.83; 0.87][-1.83;\ 0.87]
Biden Dropout $\times$ Treatment $0.36$ $1.82$
[1.53; 2.26][-1.53;\ 2.26] [0.16; 3.81][-0.16;\ 3.81]
Harris Presumptive Nomination $0.41$ $1.43^{*}$
[0.94; 1.76][-0.94;\ 1.76] [0.09; 2.78][0.09;\ 2.78]
Harris Presumptive $\times$ Treatment $0.99$ $2.19^{*}$
[0.87; 2.84][-0.87;\ 2.84] [4.12;0.26][-4.12;\ -0.26]
Harris Official Nomination 0.22-0.22 $2.96^{*}$
[1.71; 1.27][-1.71;\ 1.27] [4.67;1.24][-4.67;\ -1.24]
Harris Official Nomination $\times$ Treatment $0.92$ $2.81^{*}$
[1.24; 3.09][-1.24;\ 3.09] [0.28; 5.34][0.28;\ 5.34]
Num. obs. $6,116$ $4,778$
Num. groups $622$ $468$
Null hypothesis value outside the confidence interval.
Table S32: Affective Polarization: In-Feed. Regression tables of the analysis of the affective polarization outcome for the in-feed survey responses, controlling for major political events.